# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 1.2.1. The non-contraction property
Let $G=\mathbf{G}(\mathbb{R})$ be such that the identity component $G^{0}$ is generated by unipotent one-parameter subgroups. Suppose $\rho:G\to\text{SL}(n,\mathbb{R})$ is rational homomorphism with a finite kernel. Then, for any polynomially bounded definable curve $\varphi:[0,\infty)\to G$, if $\rho\circ\varphi$ is non-contracting in $\text{SL}(n,\mathbb{R})$, then $\varphi$ is non-contracting in $G$.
1.3. The main equidistribution result
In Section 2.2, we will show that for a given continuous definable non-contracting curve $\varphi:[0,\infty)\to G$ in an algebraic group $G=\mathbf{G}(\mathbb{R})$, there exists a unique subgroup generated by one-parameter unipotent subgroups of $G$, denoted by $H_{\varphi}\leq G$ and called the hull of $\varphi$, such that $\varphi([0,\infty))/H_{\varphi}$ is contained in a compact subset of $G/H_{\varphi}$. Also, we can choose a bounded definable curve $\beta:[0,\infty)\to G$, which we refer to as a correcting curve such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t$; see Theorem 2.4 and Definition 2.5. We note that $\lim_{t\to\infty}\beta(t)=\beta_{\infty}\in G$ by the Monotonicity Theorem [Geom_cat_and_o_min_str, Theorem 4.1].
Theorem 1.8.
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\mathcal{G}\leq G$ is a connected, closed Lie subgroup with a lattice $\Gamma\leq\mathcal{G}$. Suppose that $\varphi:[0,\infty)\to G$ is a continuous, unbounded, non-contracting curve in $G$ which is definable in a polynomially bounded o-minimal structure. Suppose that $\varphi\left([0,\infty)\right)\subset\mathcal{G}$. Let $H_{\varphi}$ be the hull of $\varphi$ in $G$ and let $\beta$ be a definable correcting curve in $G$, so that $\beta\varphi\subset H_{\varphi}$, see Definition 2.5. Then $H_{\varphi}\subset\mathcal{G}$, $\beta([0,\infty))\subset\mathcal{G}$, and the following holds: Fix $x_{0}\in\mathcal{G}/\Gamma$. Then,
$\lim_{T\to\infty}\mu_{T,\beta\varphi,x_{0}}=\mu_{\overline{H_{\varphi}x_{0}}},$
in the weak-$\ast$ topology, where, in view of Ratner’s theorem, $\mu_{\overline{H_{\varphi}x_{0}}}$ is the $L$-invariant probability measure on $\overline{H_{\varphi}x_{0}}=Lx_{0}$ for a closed connected subgroup $L$ of $\mathcal{G}$. In particular,
$\lim_{T\to\infty}\mu_{T,\varphi,x_{0}}=\beta_{\infty}^{-1}\mu_{\overline{H_{\varphi}x_{0}}},$
where $\beta_{\infty}=\lim_{t\to\infty}\beta(t)$.
Remark 1.9.
the non-contraction assumption is necessary for Theorem 1.8 to hold in the above generality. We note that if $\varphi$ is a contracting curve in $\text{SL}(n,\mathbb{R})$, then there exists $x\in\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that $\lim_{t\to\infty}\varphi(t)x=\infty$. To see the existence of such $x$, since $\varphi$ is contracting, by Proposition 1.6, there exists linearly independent vectors $\{v_{1},...,v_{k}\}\subseteq\mathbb{R}^{n}$ such that $\varphi(t).(v_{1}\wedge\cdots\wedge v_{k})\to 0$ as $t\to\infty$. Identify SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ with the space of unimodular lattices, and let $x$ be a lattice which includes $\{v_{1},...,v_{k}\}$. Then by Minkowski’s theorem, the length of the shortest nonzero vector(s) of the discrete subgroup $\text{Span}_{\mathbb{Z}}\{\varphi(t)v_{1},...,\varphi(t)v_{k}\}\subseteq\varphi(t)x_{0}$ goes to zero as $t\to\infty$. By Mahler’s criterion, the curve $\varphi(t)x_{0}$ diverges to $\infty$ in the space of unimodular lattices.
In the next result, proved in Section 7, we give an example of a definable curve in $\text{SL}(n+1,\mathbb{R})$ whose trajectory diverges in every irreducible representation of $\text{SL}(n+1,\mathbb{R})$; in other words, it is non-contracting, and its hull is $\text{SL}(n+1,\mathbb{R})$. As a consequence, every trajectory of such a curve in the space of unimodular lattice in $\mathbb{Z}^{n+1}$ gets equidistributed (Corollary 1.11).
Proposition 1.10.
Let $f_{0},...,f_{n}$, and $h_{1},\ldots,h_{n}$ be continuous real functions on $[0,\infty)$ definable in a polynomially bounded o-minimal structure. For $t\geq 0$, define
@@ 1.3. The main equidistribution result
    For any $i\in\{1,\ldots,n\}$, $\deg(f_{i}+g)>n-i$ for any $g$ in the liner span of $f_{i+1},\ldots,f_{n}$.
@@
Let $G$ be a Lie group and $\Gamma$ be a lattice in $G$. Suppose that $\rho:\text{SL}(n+1,\mathbb{R})\to G$ is a continuous homomorphism. Let $\varphi:[0,\infty)\to\text{SL}(n+1,\mathbb{R})$ be a curve satisfying the conditions of Proposition 1.10. Then for any $x\in G/\Gamma$, and any bounded continuous function $f$ on $G/\Gamma$,
@@ 1.4. Growth of o-minimal functions
We now describe the $(C,\alpha)$-good property concerning the growth of certain families of functions definable in polyonomially bounded o-minimal structures. The property is well known for polynomials, and it is new in the o-minimal setting. It is key for the technique used in this paper.
We need the following definition.
@@
Let $\mathcal{V}$ be finite dimensional vector space spanned over $\mathbb{R}$ by real functions definable in a polynomially bounded o-minimal structure. Identify $\mathcal{V}$ with $\mathbb{R}^{N}$ by choosing a basis for $\mathcal{V}$. A subset $\mathscr{F}\subseteq\mathcal{V}$ is called a closed definable cone if it is a closed definable subset of $\mathbb{R}^{N}$ such that if $v\in\mathscr{F}$ then $\alpha v\in\mathscr{F}$ for all scalars $\alpha\geq 0$.
Example 1.13.
[…]
$\mathcal{V}:=\text{Span}_{\mathbb{R}}\{\|\psi(t)\textbf{v}\|^{2}:\textbf{v}\in\mathbb{R}^{m}\}$
is finite dimensional, and :=$\mathscr{F}:=\{\|\psi(t)\textbf{v}\|^{2}:\textbf{v}\in\mathbb{R}^{m}\}$ is a closed definable cone.
@@
($(C,\alpha)$-good property). Let $\mathcal{V}$ be a finite-dimensional real vector space spanned by functions $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure. Suppose that $\mathscr{F}\subset\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\smallsetminus\{0\}$ it holds that
@@
Then, there exist $C>0$, $\alpha>0$ and $T_{0}\geq 1$ such that for any $f\in\mathscr{F}\smallsetminus\{0\}$, $f$ is $(C,\alpha)$-good in $[0,\infty)$; that is, $\forall\epsilon>0$, and for every bounded interval $I\subseteq[T_{0},\infty)$
We note the following result is an immediate consequence of Theorem 1.14.
@@ 1.5. Proof ideas and structure of the paper
Our main theorem will be proved by the following strategy which is common in many works on equidistribution in homogeneous spaces, see e.g. [Dani1993LimitDO, Mozes1995OnTS]:
@@
    Proving that there is no escape of mass – that is, any limiting measure $\mu$ of the probability measures $\mu_{T}$ is a probability measure.
@@
    Showing that any limiting measure $\mu$ of the measures $\mu_{T}$ is invariant under a group generated by unipotent one-parameter subgroups.
@@
    Studying the ergodic decomposition of a limiting measure $\mu$. Since $\mu$ is invariant under a unipotent group, every ergodic component appearing in the ergodic decomposition of $\mu$ is homogeneous by Ratner’s measure classification theorem. The theorem is proved once it is verified that $\mu$ equals to exactly one uniquely deteremined ergodic component. This will step will be treated using the linearization technique of [Dani1993LimitDO].
@@
The $(C,\alpha)$-good property will be used to prove the non-escape of mass (Section 4) and to apply the linearization technique for the ergodic decomposition of our limiting measure (Section 6).
1.6. Some conventions about notation
We will denote the Zariski closure of a subset $A$ of an affine space by $\text{zcl}(A)$. For a linear Lie group $H$, we denote by $H_{u}$ the subgroup of $H$ generated by all one-parameter unipotent subgroups contained in $H$. We will write $G:={\bf G}(\mathbb{R})$ to mean that $G$ is the set of $\mathbb{R}$ points of an algebraic group ${\bf G}$ defined over $\mathbb{R}$.
@@ Acknowledgement.
We thank Chris Miller for helpful discussions.
2. Consequences of the non-contracting property
@@ 2.1. The criterion for the non-contracting property
Our goal here is to prove Proposition 1.6, namely to reduce the verification of the non-contracting property for any vector in any representation to decomposable vectors in the exterior of the standard representation of $\text{SL}(n,\mathbb{R})$. Our proof follows from the observations in [shah2023equidistribution, Section 2], which uses Kempf’s numerical criteria on geometric invariant theory [Kem78].
We recall the following preliminaries. Let S denote the full diagonal subgroup of $\text{SL}(n,\mathbb{R})$. The group of strictly upper triangular matrices in $\text{SL}(n,\mathbb{R})$, denoted by $N$, is a maximal unipotent subgroup of $\text{SL}(n,\mathbb{R})$ normalized by S. Let $X^{\ast}(\textbf{S})$ denote the group of algebraic homomorphisms from S to $\mathbb{R}^{\ast}$ defined over $\mathbb{R}$. Then $X^{\ast}(\textbf{S})$ is a free abelian group on $(n-1)$-generators, and we treat it as an additive group. For each $i\in\{1,\ldots,n-1\}$, let $\alpha_{i}\in X^{\ast}(\textbf{S})$ be defined by $\alpha_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{i}/t_{i+1}$. Then $\Delta=\{\alpha_{i}:1\leq i\leq n-1\}$ is the set of simple roots on S corresponding to the choice of the maximal unipotent subgroup $N$. For each $11\leq i\leq n-$ 1, let $\mu_{i}\in X^{\ast}(\textbf{S})$ be defined by $\mu_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{1}\cdots t_{i}$. Then $\{\mu_{i}:1\leq i\leq n-1\}$ is the set of fundamental characters (weights) of $\mathbf{S}$ with respect to our choice of the simple roots. For each $11\leq i\leq n-$ 1, let $W_{i}$ denote the $i$-th exterior of the standard representation of $\text{SL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$, called a fundamental representation of $\text{SL}(n,\mathbb{R})$. So $W_{i}=\wedge^{i}\mathbb{R}^{n}$, and for any $v_{1}\wedge\cdots\wedge v_{i}\in W_{i}$ and $g\in\text{SL}(n,\mathbb{R})$, $g.(v_{1}\wedge\cdots\wedge v_{i})=gv_{1}\wedge\cdots\wedge gv_{i}$. Let $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$. Then $w_{i}$ is fixed by $N$, and S acts on the line $\mathbb{R}w_{i}$ via the fundamental character $\mu_{i}$.
A dominant integral character (weight) is a non-negative integral combination of fundamental characters. Suppose $\chi=m_{1}\mu_{1}+\cdots+m_{n-1}\mu_{n-1}$, where $m_{i}\in\mathbb{Z}_{\geq 0}$. For each $i$, let $W_{i}^{\otimes m_{i}}$ denote the tensor product of $m_{i}$-copies of $W_{i}$, and $w_{i}^{m_{i}}:=w_{i}\otimes\cdots\otimes w_{i}\in W_{i}^{\otimes m_{i}}$. Consider the tensor product representation of $\text{SL}(n,\mathbb{R})$ on $W_{\chi}=W_{1}^{\otimes m_{1}}\otimes\cdots\otimes W_{n-1}^{\otimes m_{n-1}}$, and let $w_{\chi}=w_{1}^{m_{1}}\otimes\cdots\otimes w_{n-1}^{m_{n-1}}$. Then $w_{\chi}$ is fixed by $N$, and S acts on the line $\mathbb{R}w_{\chi}$ via the character $\chi$.
@@
Let $V$ be a representation of $G:=\text{SL}(n,\mathbb{R})$. Suppose $v\in V$ is a a nonzero such that $\varphi(t)v\to 0$ as $t\to\infty$. We call such a $v$ a $G$-unstable vector in $V$. So, by [shah2023equidistribution, Remark 2.3], there exist $g_{0}\in G$, a dominant integral character $\chi\in X^{\ast}(\textbf{S})$, and constants $\beta>0$ and $C>0$ such that for any $g\in G$,
(2.1) $\lVert gg_{0}.w_{\chi}\rVert\leq C\lVert g.v\rVert^{\beta}.$
Therefore, $\lim_{t\to\infty}\varphi(t)g_{0}.w_{\chi}=0$. Let $A=\textbf{S}^{0}$, the component of the identity in S. Let $K=\text{SO}(n)$. Then by Iwasawa decomposition, $G=KAN$. For each $t$, we express $\varphi(t)g_{0}=k_{t}a_{t}n_{t}$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$. Without loss of generality, we may assume that the norm on $W_{\chi}$ is $K$-invariant. Since, $N$ fixes $w_{\chi}$,
$\lVert\varphi(t)g_{0}.w_{\chi}\rVert=\lVert a_{t}.w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
Hence, since $\lim_{t\to\infty}\varphi(t)g_{0}.w_{\chi}=0$, we get $\lim_{t\to\infty}\chi(a_{t})=0$. Now
@@
$\lVert\varphi(t_{l})g_{0}.w_{i}\rVert=\lVert{a_{t_{l}}.w_{i}}\rVert=\mu_{i}(a_{t_{l}})\lVert w_{i}\rVert\to 0\text{, as $l\to\infty$.}$
Since the coordinate functions of $t\to\varphi(t)g_{0}.w_{i}$ are definable in an o-minimal structure, we get $\lim_{t\to\infty}\varphi(t)g_{0}.w_{i}=0$. Let $v_{j}=g_{0}e_{j}$ for each $j$. Since $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$, $\varphi(t).(g_{0}w_{i})=\varphi(t).(v_{1}\wedge\cdots\wedge v_{i})\to 0$ as $t\to\infty$. This contradicts our assumption 1.4. ∎
The following statement will be used in the proof of Lemma 7.1, which will needed for proving Propositions 1.10.
@@
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve such that
(2.2) $\lim_{t\to\infty}\varphi(t).(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$. Then for every finite dimensional rational representation $V$ and an $v\in V$ such that the curve $\varphi(t).v,t\geq 0$ is bounded in $V$, it holds that $\text{SL}(n,\mathbb{R}).v$ is Zariski closed.
@@
We will use the notations above. Suppose that $V$ is a finite dimensional rational representation of $\text{SL}(n,\mathbb{R})$ and let $v\in V$ such that $\text{SL}(n,\mathbb{R}).v$ is not Zariski closed. By [shah2023equidistribution, Corollary 2.5] and by [shah2023equidistribution, Remark 2.3], there exist a $g_{0}\in\text{SL}(n,\mathbb{R})$, a dominant integral character $\chi\in X^{\ast}(\textbf{S})$, and a constant $\beta>0$ such that for any $R>0$ there is a constant $C>0$ with the following property:
(2.3) $\text{if $\|g.v\|\leq R$, then }\lVert gg_{0}.w_{\chi}\rVert\leq C\lVert g.v\rVert^{\beta}.$
Now suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve satisfying the diverging property (2.2) and assume for contradiction that $\varphi(t).v,~{}t\geq 0$ is bounded in $V$. Then, by (2.3), we get that $\lVert\varphi(t)g_{0}.w_{\chi}\rVert,~{}t>0$ is bounded. As in the proof of Proposition 1.6, we use Iwasawa decomposition to express $\varphi(t)g_{0}=k_{t}a_{t}n_{t},t>0$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$, and observe that
$\lVert\varphi(t)g_{0}.w_{\chi}\rVert=\lVert a_{t}.w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
@@
for $m_{i}\geq 0$, and
$\lVert\varphi(t)g_{0}.w_{i}\rVert=\lVert{a_{t}.w_{i}}\rVert=\mu_{i}(a_{t})\lVert w_{i}\rVert,$
where $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$ and $11\leq i\leq n-$ 1. Finally, due to (2.2), we have $\lim_{t\to\infty}\lVert\varphi(t)g_{0}.w_{i}\rVert=\infty.$

### v2

@@ 1.2. The main equidistribution result
Let $f_{0},...,f_{n}$, and $h_{1},\ldots,h_{n}$ be continuous real valued functions on $[0,\infty)$ definable in a polynomially bounded o-minimal structure. For $t\geq 0$, define
@@
    For any $i\in\{1,\ldots,n\}$, $\deg(f_{i}+g)>n-i$ for any $g$ in the linear span of $f_{i+1},\ldots,f_{n}$.
@@
Let $\mathcal{G}$ be a Lie group and $\Gamma$ be a lattice in $\mathcal{G}$. Suppose that $\rho:\text{SL}(n+1,\mathbb{R})\to\mathcal{G}$ is a continuous homomorphism. Let $\varphi:[0,\infty)\to\text{SL}(n+1,\mathbb{R})$ be a curve satisfying the conditions of Proposition 1.9. Then, for any $x\in\mathcal{G}/\Gamma$, and any bounded continuous function $f$ on $\mathcal{G}/\Gamma$,
@@ 1.3. Growth of o-minimal functions
We now describe the $(C,\alpha)$-good property, introduced by Kleibock and Margulis [Klein_Marg_flows_and_dioph], concerning the growth of certain families of functions definable in polynomially bounded o-minimal structures. This property is well known for polynomials [Dani1993LimitDO, 4.1 Lemma], but is new in the o-minimal setting. It is crucial for the techniques used in this paper.
@@
Let $C$ and $\alpha$ be positive constants. We say that a function $f:[0,\infty)\to\mathbb{R}$ is $(C,\alpha)$-good on $[T_{0},\infty)\subseteq[0,\infty)$, if for every $\epsilon>0$, and for every bounded interval $I\subseteq[T_{0},\infty)$, it holds that:
For the following statement, we say that a subset $\mathscr{F}$ of a finite-dimensional vector space $\mathcal{V}$ over $\mathbb{R}$ is a definable cone, if $\mathscr{F}$ is a cone in the vector space, and $\mathscr{F}$ is a definable set when identifying $\mathcal{V}\cong\mathbb{R}^{n}$ by choosing a basis.
@@
($(C,\alpha)$-good property). Let $\mathcal{V}$ be a finite-dimensional real vector space spanned by functions $f:[0,\infty)\to\mathbb{R}$, each definable in the same polynomially bounded o-minimal structure. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\setminus\{0\}$ it holds that
@@
Then, there exist $C>0$, $\alpha>0$ and $T_{0}\geq 1$ such that for any $f\in\mathscr{F}\setminus\{0\}$, $f$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
We obtain the following statement as a direct corollary of Theorem 1.12.
@@
[…]
$\mathcal{V}:=\text{Span}_{\mathbb{R}}\{t\mapsto\|\psi(t)\textbf{v}\|^{2},\ t\geq 0:\textbf{v}\in\mathbb{R}^{m}\}$
is finite dimensional, and :=$\mathscr{F}:=\{t\mapsto\|\psi(t)\textbf{v}\|^{2},\ t\geq 0:\textbf{v}\in\mathbb{R}^{m}\}$ is a closed definable cone.
@@ 1.4. Proof ideas and structure of the paper
Our main theorem will be proved by the following strategy employed in many works on equidistribution in homogeneous spaces, for example [Dani1993LimitDO, Mozes1995OnTS, Shah1994LimitDO]:
@@
    Proving that there is no escape of mass (Section 4) – that is, any limiting measure $\mu$ of the probability measures $\mu_{T,\varphi,x}$ is a probability measure.
@@
    Showing that any limiting measure $\mu$ of the measures $\mu_{T,\varphi,x}$ is invariant under a group generated by nontrivial unipotent one-parameter subgroups (Section 5).
@@
    Studying the ergodic decomposition of a limiting measure $\mu$ (Section 6). By Ratner’s theorem [Ratner91a], every ergodic component appearing in the ergodic decomposition of $\mu$ is homogeneous. The theorem is proved once it is verified that $\mu$ is equal to exactly one uniquely determined ergodic component. For this step, we will use the linearization technique of [Dani1993LimitDO] to prove avoidance of singular sets that contain ergodic components of various types.
@@
The $(C,\alpha)$-good property will be used to prove the non-escape of mass, which is avoidance of the point at infinity, and to apply the linearization technique.
1.5. Notation and conventions
We will denote the Zariski closure of a subset $A$ of $\mathbb{R}^{N}$ by $\text{zcl}(A)$. For a linear Lie group $H$, we denote by $H_{u}$ the subgroup of $H$ generated by all the one-parameter unipotent subgroups contained in $H$. We note that if $H\subseteq\text{GL}(n,\mathbb{R})$, then $H_{u}$ equals the connected component of the identity in $\text{zcl}(H_{u})$.
@@ Acknowledgement.
We thank Chris Miller for helpful discussions. We thank anonymous referees for their careful reading and valuable suggestions which helped us improve the readability of this article.
2. On the non-contracting property and the hull
This section is divided into two main parts. In the first part, we prove Proposition 1.6, which gives a convenient criterion for verifying the non-contraction property. The second part of the section concerns the existence of the hull and the correcting curve, and discusses their behavior under homomorphisms.
@@ 2.1. The criterion for the non-contracting property
Our main goal here is to prove Proposition 1.6, namely to reduce the verification of the non-contracting property for any vector in any representation to decomposable vectors in the exterior of the standard representation of $\text{SL}(n,\mathbb{R})$. Our proof follows from the observations in [shah2023equidistribution, Section 2], which uses Kempf’s numerical criteria on geometric invariant theory [Kem78].
We recall the following preliminaries. Let ${S}$ denote the full diagonal subgroup of $\text{SL}(n,\mathbb{R})$. The group of the unipotent upper triangular matrices in $\text{SL}(n,\mathbb{R})$, denoted by $N$, is a maximal unipotent subgroup of $\text{SL}(n,\mathbb{R})$ normalized by ${S}$. Let $X^{\ast}({S})$ denote the group of algebraic homomorphisms from ${S}$ to $\mathbb{R}^{\ast}$ defined over $\mathbb{R}$. Then $X^{\ast}({S})$ is a free abelian group with $(n-1)$ generators, and we treat it as an additive group. For each $i\in\{1,\ldots,n-1\}$, let $\alpha_{i}\in X^{\ast}({S})$ be defined by $\alpha_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{i}/t_{i+1}$. Then $\Delta=\{\alpha_{i}:1\leq i\leq n-1\}$ is the set of simple roots on ${S}$ corresponding to the choice of the maximal unipotent subgroup $N$. For each $11\leq i\leq n-$ 1, let $\mu_{i}\in X^{\ast}({S})$ be defined by $\mu_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{1}\cdots t_{i}$. Then $\{\mu_{i}:1\leq i\leq n-1\}$ is the set of fundamental characters (weights) of ${S}$ with respect to our choice of simple roots. For each $11\leq i\leq n-$ 1, let $W_{i}$ denote the $i$-th exterior of the standard representation of $\text{SL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$, called a fundamental representation of $\text{SL}(n,\mathbb{R})$. So $W_{i}=\wedge^{i}\mathbb{R}^{n}$, and for any $v_{1}\wedge\cdots\wedge v_{i}\in W_{i}$ and $g\in\text{SL}(n,\mathbb{R})$, $g\cdot(v_{1}\wedge\cdots\wedge v_{i})=gv_{1}\wedge\cdots\wedge gv_{i}$. Let $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$. Then $w_{i}$ is fixed by $N$, and ${S}$ acts on the line $\mathbb{R}w_{i}$ via the fundamental character $\mu_{i}$.
A dominant integral character (weight) is a non-negative integral combination of fundamental characters. Suppose $\chi=m_{1}\mu_{1}+\cdots+m_{n-1}\mu_{n-1}$, where $m_{i}\in\mathbb{Z}_{\geq 0}$. For each $i$, let $W_{i}^{\otimes m_{i}}$ denote the tensor product of $m_{i}$-copies of $W_{i}$, and $w_{i}^{m_{i}}:=w_{i}\otimes\cdots\otimes w_{i}\in W_{i}^{\otimes m_{i}}$. Consider the tensor product representation of $\text{SL}(n,\mathbb{R})$ on $W_{\chi}=W_{1}^{\otimes m_{1}}\otimes\cdots\otimes W_{n-1}^{\otimes m_{n-1}}$, and let $w_{\chi}=w_{1}^{m_{1}}\otimes\cdots\otimes w_{n-1}^{m_{n-1}}$. Then $w_{\chi}$ is fixed by $N$, and ${S}$ acts on the line $\mathbb{R}w_{\chi}$ via the character $\chi$. We endow $W_{\chi}$ with a $\text{SO}(n,\mathbb{R})$-invariant norm.
@@
Let $V$ be a finite dimensional representation of $G:=\text{SL}(n,\mathbb{R})$ endowed with some norm. Suppose $v\in V$ is a nonzero vector such that $\varphi(t)\cdot v\to 0$ as $t\to\infty$. Then $\overline{G\cdot v}\ni 0$, so $v$ is a $G$-unstable vector in $V$. So, by [shah2023equidistribution, Remark 2.3], there exist $g_{0}\in G$, a dominant integral character $\chi\in X^{\ast}({S})$, and constants $\beta>0$ and $C>0$ such that for any $g\in G$,
(2.1) $\lVert gg_{0}\cdot w_{\chi}\rVert\leq C\lVert g\cdot v\rVert^{\beta}.$
Therefore, $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{\chi}=0$. Let $A={S}^{0}$, the component of the identity in ${S}$. Let $K=\text{SO}(n,\mathbb{R})$. Then by Iwasawa decomposition, $G=KAN$. For each $t$, we express $\varphi(t)g_{0}=k_{t}a_{t}n_{t}$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$. Since the norm on $W_{\chi}$ is $K$-invariant and $N$ fixes $w_{\chi}$,
$\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert=\lVert a_{t}\cdot w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
By (2.1), $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{\chi}=0$. Therefore, $\lim_{t\to\infty}\chi(a_{t})=0$. Now
@@
$\lVert\varphi(t_{l})g_{0}\cdot w_{i}\rVert=\lVert{a_{t_{l}}\cdot w_{i}}\rVert=\mu_{i}(a_{t_{l}})\lVert w_{i}\rVert\to 0\text{, as $l\to\infty$.}$
Since the coordinate functions of $t\to\varphi(t)g_{0}\cdot w_{i}$ are definable in an o-minimal structure, we get $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{i}=0$. Let $v_{j}=g_{0}e_{j}$ for each $j$. Since $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$, $\varphi(t)\cdot(g_{0}\cdot w_{i})=\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{i})\to 0$ as $t\to\infty$. This contradicts the assumption that the curve is non-contracting. ∎
We record the following statement. It will be used in the proof of Lemma 7.1, which is in turn needed for proving Proposition 1.9.
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a map such that
(2.2) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$. Then for any finite-dimensional rational representation $V$ and any $v\in V$, if $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$, then $\text{SL}(n,\mathbb{R})\cdot v$ is Zariski closed.
@@
Suppose that $V$ is a finite-dimensional rational representation of $G=\text{SL}(n,\mathbb{R})$ and let $v\in V$ such that $G\cdot v$ is not Zariski closed. So there exists a finite-dimensional rational representation $W$ of $G$ and a $G$-equivariant polynomial map $P:V\to W$ such that $P(v)\neq 0$ and $\overline{G\cdot P(v)}\ni 0$, see Lemma A.1. Therefore, by [shah2023equidistribution, Remark 2.3], there exist a $g_{0}\in G$, a nontrivial dominant integral character $\chi\in X^{\ast}({S})$, and constants $\beta>0$ and $C>0$ such that for any $g\in G$,
$\lVert gg_{0}\cdot w_{\chi}\rVert\leq C\lVert g\cdot P(v)\rVert^{\beta}=C\lVert P(g\cdot v)\rVert^{\beta},$
cf. [shah2023equidistribution, Corollary 2.5]. Since $P$ is continuous, given any $R>0$, there exists $C_{R}>0$ such that for any $g\in G$,
(2.3) if gvR, then $\text{if $\|g\cdot v\|\leq R$, then }\lVert gg_{0}\cdot w_{\chi}\rVert\leq C_{R}\lVert g\cdot v\rVert^{\beta}.$
Now suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a map satisfying the diverging property (2.2) and assume for contradiction that $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$. Then, by (2.3), we get that $\{\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert:t>0\}$ is bounded. As in the proof of Proposition 1.6, by Iwasawa decomposition, for all $t>0$, $\varphi(t)g_{0}=k_{t}a_{t}n_{t}$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$, and
$\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert=\lVert a_{t}\cdot w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
@@
for some $(m_{1},\ldots,m_{n-1})\in\mathbb{Z}_{\geq 0}$ and $m_{i}>0$ for some $i$. Now
$\lVert\varphi(t)g_{0}\cdot w_{i}\rVert=\lVert{a_{t}\cdot w_{i}}\rVert=\mu_{i}(a_{t})\lVert w_{i}\rVert,$
where $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$ and $11\leq i\leq n-$ 1. And by (2.2), we have $\lim_{t\to\infty}\lVert\varphi(t)g_{0}\cdot w_{i}\rVert=\infty$. So, $\lim_{t\to\infty}\mu_{i}(a_{t})\to\infty$. ∎